The length of a rectangle is 8 meters less than 5 times the width. If the perimeter of the rectangle is 104 meters find the width of the rectangle.

Answer by addingup(3677)

L = 5W-8 we'll use this value for L next
2L+2W = 104 Substitute
2(5W-8)+2W = 104
10W-16+2W = 104
12W = 120
W = 10
and
L = 5W-8 = 5(10)-8 = 42
----------------------
Check:
2(42)+2(10) = 104
84+20 = 104 Correct
:
John